### v1

# Prompt and Parameter Co-Optimization for Large Language Models

Abstract
Prompt optimization and fine-tuning are two major approaches to improve the performance of Large Language Models (LLMs). They enhance the capabilities of LLMs from complementary perspectives: the former through explicit natural language, and the latter through implicit parameter updates. However, prior work has typically studied them in isolation, leaving their synergistic potential largely underexplored. To bridge this gap, in this paper, we introduce MetaTuner, a novel framework that jointly integrates prompt optimization and fine-tuning for LLM training. Specifically, we introduce two neural networks to generate prompts and parameters, respectively, while allowing them to share a common bottom encoding layer to enable knowledge sharing. By the guidance of the final supervised signals, our framework is optimized to discover the optimal combinations between the prompts and parameters. Given that prompt learning involves discrete optimization while fine-tuning operates in a continuous parameter space, we design a supervised regularization loss to train our framework effectively. Extensive experiments across diverse benchmarks show that our method consistently outperforms the baselines.
1 Introduction
Large language models (LLMs) have demonstrated remarkable success across a wide range of domains, including e-commerce (ref:shoppingmmlu, ; ref:ecellm, ; ref:ecomgpt, ; ref:ecommerce, ), education (ref:edu1, ; ref:edu2, ; ref:edu3, ), and social science (ref:social1, ; ref:social2, ; ref:social3, ; ref:social4, ). Unlike traditional machine learning models, whose capabilities primarily stem from parameter tuning, LLMs can enhance their performance through two complementary strategies: prompt optimization and fine-tuning.
Previously, prompt optimization and fine-tuning are often treated as independent research lines. For example, in the line of prompt optimization, current methods (ref:hmaw, ; ref:grad_sum, ; ref:llm_dual_phase_po, ; ref:rlprompt_improve, ) mainly focus on discovering or learning appropriate input context that can elicit the desired behavior from frozen language models. One group of works (e.g., OPRO (ref:opro, )) focuses on optimizing prompts with LLMs, whereas another group (e.g., RLPrompt (ref:rlprompt, ), BPO (ref:bpo, )) trains an auxiliary model for prompt generation. In the line of fine-tuning, existing approaches concentrate on better training algorithms to adjust the model parameters. Supervised Fine-Tuning (SFT) proposes to optimize the model with cross-entropy loss, while more recently, RL-based approaches such as RLHF (ref:rlhf, ) and DPO (ref:dpo, ) are developed to optimize the model based on human feedback signals.
While the above models have shown promising results in various domains, both prompt optimization and fine-tuning have their own limitations. Specifically, prompt optimization typically focuses on refining general instructions within prompts to elicit the capabilities of LLMs. However, it often falls short in adapting the pretrained parameters to the complex patterns present in large-scale, task-specific datasets, especially when the knowledge encoded in the prompts and the model parameters may conflict. Fine-tuning, on the other hand, usually leverages manually crafted prompts as inputs during model training. However, the choice of input prompts can significantly affect fine-tuning outcomes, and existing methods offer no guarantee that these prompts are optimal for maximizing downstream performance. To further validate these statements, we conduct a series of preliminary experiments, and the corresponding results are presented in Figure 1(a). As illustrated in the left subfigure, by training the LLM parameters based on the datasets of MATH (ref:math, ) and HotpotQA (ref:hotpotqa, ), the fine-tuning methods substantially outperform the prompt optimization strategies on average. However, the effectiveness of fine-tuning is highly sensitive to the choice of prompts. As shown in the right subfigure, when sub-optimal prompts are used for training, the performance degrades and can even fall below that of prompt optimization methods. Motivated by the above mechanistic and empirical analyses, a natural question arises: Can we design a unified framework to combine prompt optimization and fine-tuning to mitigate the weaknesses of each other?
(a) Preliminary analysis of prompt optimization and fine-tuning methods. In the left subfigure, we present a comparison between four representative prompt optimization strategies and four popular fine-tuning methods. In the right subfigure, we utilize SFT as the backbone fine-tuning method and further evaluate the model with different training prompts, where the corresponding results are marked with stars. Additional details for the implementation can be found in the Appendix D.1.
Answering the above question is not straightforward. To begin with, prompt optimization and fine-tuning have different execution processes and optimization objectives. The former optimizes the model externally by finding the suitable input context to activate existing knowledge, while the latter modifies the model internally by updating parameters to better fit the target data distribution. Effectively reconciling these differences to combine the two approaches requires careful design. In addition, prompt optimization is inherently a discrete optimization problem, whereas fine-tuning operates in a continuous optimization space. Combining the two, therefore, introduces mixed optimization challenges, such as non-differentiable gradients. Effectively leveraging the final supervision signal to address this discrete-continuous optimization problem remains a significant challenge.
To overcome the above challenges, we propose MetaTuner, a novel framework for joint prompt optimization and fine-tuning. Our approach first employs a meta encoder to encode the input query. The resulting embedding is then fed into two separate decoders—a prompt decoder and a parameter decoder—which simultaneously generate query-specific prompts and parameters for downstream large language models. This design allows the prompts and parameters to share the common knowledge from the meta encoder, promoting their mutual improvement. Considering that prompt learning is a discrete process while the parameter training is continuous, we design a supervised regularization loss to solve the mixed discrete-continuous optimization problem. Specifically, we first select the optimal rollout prompts to optimize the prompt decoder in a supervised learning manner, and then share the parameters of the prompt and parameter decoders for joint differentiable training.
The main contributions of this paper are summarized as follows: (1) We propose the idea of jointly optimizing the prompts and parameters of LLMs to leverage their complementary strengths and mitigate each other’s weaknesses. (2) To realize this idea, we introduce a framework called MetaTuner, which unifies the prompt and parameter generation processes to enhance the overall task performance. (3) We design a supervised regularization objective to address the non-differentiability challenge and effectively optimize the MetaTuner framework. (4) We conduct extensive experiments on MATH, GSM8K, HotpotQA, and CosmosQA to demonstrate the effectiveness and robustness of our approach.
2 Problem Formulation
In LLMs, we typically work with an input-output pair $(x,y)$, where $x$ represents a question such as “What is the capital of France?”, and $y$ corresponds to the expected answer like “Paris”. Unlike traditional machine learning models, LLMs often require an additional prompt $p$ to rephrase the input $x$ in a way that better guides the model to generate the desired output $y$. In the example above, $p$ could be “Answer the following question briefly:”. Letting the language model be denoted as $\mathcal{M}$, the overall process can be formulated as:
${y}=\mathcal{M}(p,x),$ (1)
To enhance the capability of $\mathcal{M}$, there are two primary research directions: (1) Prompt optimization. The core idea of this approach is to find an optimal prompt $p$ that facilitates more accurate prediction of $y$. Given a loss function $\mathcal{L}$, the objective can be formulated as:
$\min_{p_{i}}\sum_{i=1}^{N}\mathcal{L}(\mathcal{M}(p_{i},x_{i}),y_{i}),$ (2)
where $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ denotes the set of input-output pairs. We use the subscript $i$ in $p_{i}$ to indicate that the prompt is input-sensitive. However, depending on the specific application, a shared or unified prompt across all inputs is also a natural extension. (2) Fine-tuning. This method regards the LLM as a traditional machine learning model, and optimizes its parameters based on the given input-output pairs. In specific, the objective is formulated as follows:
$\min_{\theta}\sum_{i=1}^{N}\mathcal{L}(\mathcal{M}_{\theta}(p_{i},x_{i}),y_{i}),$ (3)
where we use $\theta$ to represent the parameters of the large language model.
Previous studies have employed various strategies (ref:rlhf, ; ref:dpo, ; ref:kto, ; ref:opro, ; ref:bpo, ; ref:rlprompt, ; ref:dora, ; ref:pissa, ) to instantiate objectives (2) and (3), respectively. Here, we take a different perspective by formulating them within a unified framework, laying the foundation for letting prompt optimization and fine-tuning mutually enhance each other. Specifically, we regard the prompt as a “special parameter”, then objectives (2) and (3) can be unified by the following loss function:
$\min_{\theta,p_{i}}\sum_{i=1}^{N}\mathcal{L}(\mathcal{M}_{\theta}(p_{i},x_{i}),y_{i}).$ (4)
Compared with previous studies, this loss function considers prompts and model parameters as two complementary dimensions that jointly influence the prediction from $x_{i}$ to $y_{i}$. By optimizing them within a unified objective, we are actually finding the optimal combinations between the prompts and parameters to better fit specific tasks. The loss function can help to regularize any sub-optimal choices in either the prompt or parameter space. Readers may find that the above objective can also be applied to soft-prompting strategies (ref:p_tuning, ; ref:p_tuning_v2, ; ref:prefix_tuning, ; ref:google_soft_prompt, ; ref:prompt_tuning, ). However, in this paper, we focus on generating explicit prompts, because the natural language prompts are more explainable, transparent, and human-friendly.
3 The MetaTuner Framework
Figure 1: Illustration of the MetaTuner framework. The input query is first encoded by the meta encoder, and then two parallel decoders are utilized to generate the prompts and parameters, respectively, which are finally applied to the downstream actor model for problem solving.
The overall architecture of our framework can be seen in Figure 1. We first design a neural network with continuous parameters to generate prompts, effectively transforming the discrete optimization problem into a continuous task. Next, we unify prompt optimization and fine-tuning by introducing a shared meta encoder that simultaneously produces input prompts and model parameters, allowing mutual adaptation to better fit the training samples. Finally, we design a supervised regularization loss to address the non-differentiable problem, enabling effective and stable framework optimization.
3.1 Prompt Generation with Continuous Neural Network
In practice, a prompt typically consists of dozens or even hundreds of words. Generating such a prompt entirely from scratch, without any reference, poses a significant challenge. To mitigate this difficulty, we first identify an initial prompt—albeit potentially sub-optimal—and then leverage a large language model (LLM) to rewrite it, thereby simplifying the prompt generation task. To reduce the effort of assigning an independent initial prompt to each input $x_{i}$, we let the LLM automatically explore the prompt space, starting from the same initial prompt. Formally, let $\mathcal{G}$ denote the model for prompt generation, then the rewritten prompt is represented as follows:
${p}_{i}=\mathcal{G}_{\phi}(\tilde{p},{x}_{i}),$ (5)
where we use $\phi$ to indicate the learnable parameters of $\mathcal{G}$, and $\tilde{p}$ is the predefined initial prompt. Based on equation (5), the objective (4) for LLM training is improved as follows:
$\min_{\theta,\phi}\sum_{i=1}^{N}\mathcal{L}(\mathcal{M}_{\theta}(\mathcal{G}_{\phi}(\tilde{p},{x}_{i}),x_{i}),y_{i}),$ (6)
where the optimizable parameters $\theta$ and $\phi$ are all continuous.
3.2 Joint Prompt and Parameter Generation
Intuitively, both prompts and model parameters aim to enhance the prediction of outputs $y_{i}$ from given inputs $x_{i}$, and are therefore expected to share common underlying knowledge. Furthermore, as outlined in the introduction, prompt optimization and fine-tuning each exhibit distinct limitations. This motivates the need for a principled approach that enables the two strategies to complement one another and mitigate each other’s shortcomings.
To incorporate the above intuitions into our model design, we do not treat $\theta$ as independent parameters. Instead, we generate them using a neural network. This approach allows us to share parameters between the prompt generation and parameter generation models, thereby establishing a common foundation for both prompt optimization and fine-tuning. Formally, we partition the parameters of the prompt generation model $\mathcal{G}$ into two components: $\phi=\{\phi_{s},\phi_{p}\}$. Here, $\phi_{s}$ denotes the shared parameters, which are also utilized for LLM parameter generation, while $\phi_{p}$ refers to the parameters exclusively used for prompt generation. We denote the model for LLM parameter generation as $\mathcal{F}$, and its parameters are also separated into two parts $\{\phi_{s},\phi_{q}\}$, where $\phi_{q}$ is tailored to generating LLM parameters. Based on these formulation, we further improve objective (4) to obtain the final optimization target as follows:
$\min_{\phi_{s},\phi_{p},\phi_{q}}\sum_{i=1}^{N}\mathcal{L}(\mathcal{M}_{\mathcal{F}_{(\phi_{s},\phi_{q})}(\tilde{p},{x}_{i})}(\mathcal{G}_{(\phi_{s},\phi_{p})}(\tilde{p},{x}_{i}),x_{i}),y_{i}).$ (7)
In this target, fine-tuning and prompt optimization are connected through a shared-private parameterization mechanism. On one hand, the shared parameters $\phi_{s}$ enable mutual regularization between the two approaches—any sub-optimal solution from one can be corrected by the other under the unified loss function. On the other hand, the private parameters $\phi_{p}$ and $\phi_{q}$ provide each approach with the flexibility to independently explore and identify its own optimal solutions.
3.3 Optimization with Supervised Regularization
In the above objective (7), the outputs of the prompt generation model $\mathcal{G}$ are discrete tokens, which makes it hard to directly compute the gradient of $\phi_{p}$. To address this challenge, we further introduce a surrogate loss based on supervised regularization as follows:
$\min_{\phi_{s},\phi_{p},\phi_{q}}\sum_{(x_{i},y_{i})\in D_{1}}\mathcal{L}(\mathcal{M}_{\mathcal{F}_{(\phi_{s},\phi_{q})}(\tilde{p},{x}_{i})}(\mathcal{G}_{(\phi_{s},{\phi}^{\prime}_{p})}(\tilde{p},{x}_{i}),x_{i}),y_{i})+\sum_{(x_{i},p_{i})\in D_{2}}\alpha\cdot\mathcal{L}(\mathcal{G}_{(\phi_{s},\phi_{p})}(\tilde{p},{x}_{i}),p_{i}).$ (8)
In this objective, the first term is identical to (7), except that the private parameters used to generate the prompt (i.e., ${\phi}^{\prime}_{p}$) are kept fixed. The second term acts as a regularizer, optimizing $\phi_{s}$ and $\phi_{p}$ under a supervised learning paradigm. Here, $D_{1}$ denotes the original training dataset, while $D_{2}$ is an expert dataset in which each $p_{i}$ yields the correct $y_{i}$ given $x_{i}$. After several optimization batches, ${\phi}^{\prime}_{p}$ is updated using $\phi_{p}$. In practice, one can optimize the above two terms either separately in sequence or jointly as a single loss. We evaluate both of these implementations in our experiments.
To address the above non-differentiable problem, we also explored strategies such as Gumbel-Softmax (ref:gumbel_softmax, ), but these methods generally produced inferior performance. We speculate that our supervised regularizer is more effective because it provides direct supervision on the optimal prompts, avoiding intermediate softening steps or approximations.
3.4 Framework Specification
In the above sections, we outlined the general design principles of our framework. In this section, we turn to the implementation details, with a particular focus on the key components $\mathcal{G}$ and $\mathcal{F}$.
Specification of $\mathcal{G}$. In our framework, an LLM such as Qwen2.5-7B (ref:qwen2.5, ) is selected as the implementation of $\mathcal{G}$. In specific, we use the first $k$ decoder layers as the meta encoder (i.e., $\phi_{s}$), while the subsequent layers serve as the prompt decoder (i.e., $\phi_{p}$). We denote the hidden states by $h_{i}\in\mathbb{R}^{l\times h_{\mathcal{G}}}$, where $l$ represents the input sequence length and $h_{\mathcal{G}}$ corresponds to the hidden size.
Specification of $\mathcal{F}$. For training efficiency, we adopt LoRA (ref:lora, ) to fine-tune the downstream LLMs, with $\mathcal{F}$ responsible for generating the LoRA parameters. Specifically, let $\Delta W=\theta_{i}^{b}\cdot\theta_{i}^{a}$ denote the update matrix, where $\theta_{i}^{b}\in\mathbb{R}^{d_{\mathcal{M}}\times r_{\mathcal{M}}}$ and $\theta_{i}^{a}\in\mathbb{R}^{r_{\mathcal{M}}\times k_{\mathcal{M}}}$. $\mathcal{F}$ first share the same $\phi_{s}$ with $\mathcal{G}$ to produce $h_{i}$. After that, we derive ${\theta}^{b}_{i}$ and ${\theta}^{a}_{i}$ from the hidden states $h_{i}$ as follows:
${\theta}^{b}_{i}$ $=\text{MM}(\text{ReLU}(\text{MM}(W_{d}^{b},h_{i})),W_{u}^{b}),$ (9)
${\theta}^{a}_{i}$ $=\text{MM}(\text{ReLU}(\text{MM}(W_{d}^{a},h_{i})),W_{u}^{a}),$
where $\phi_{q}=\{W_{d}^{b}\in\mathbb{R}^{d_{\mathcal{M}}\times l},W_{u}^{b}\in\mathbb{R}^{h_{\mathcal{G}}\times r_{\mathcal{M}}},W_{d}^{a}\in\mathbb{R}^{r_{\mathcal{M}}\times l},W_{u}^{a}\in\mathbb{R}^{h_{\mathcal{G}}\times k_{\mathcal{M}}}\}$ is the parameter decoder. ReLU is the activation function, and MM denotes the matrix multiplication operation.
4 Experiments
In this section, we conduct extensive experiments to demonstrate the effectiveness of our framework, where we focus on the following research questions (RQ):
  • RQ1: Can MetaTuner achieve better performance than existing state-of-the-art methods?
  • RQ2: Do the different components of our framework contribute to the final performance?
  • RQ3: Is the shared-private co-optimization structure in our framework effective?
  • RQ4: Is the supervised regularization loss effective in enhancing model performance?
  • RQ5: Can the capabilities of our framework generalize effectively to unseen datasets?
In the following, we answer the above questions by reporting and analyzing the experiment results.
4.1 Experimental Setup
Datasets and metrics. We evaluate MetaTuner on four datasets: MATH (ref:math, ) and GSM8K (ref:gsm8k, ) for mathematical reasoning, HotpotQA (ref:hotpotqa, ) and CosmosQA (ref:cosmosqa, ) for multi-hop and commonsense question answering. Each dataset is divided into training, development, and test sets, with details of the splits provided in the Appendix D.2. We report the exact match (EM) accuracy for MATH, GSM8K, and CosmosQA, while for HotpotQA, we adopt the F1 accuracy following the implementation of (ref:copper, ; ref:retroformer, ).
Baselines. We compare MetaTuner with a series of state-of-the-art approaches, which can be classified into four categories: (1) Vanilla Methods: Qwen2.5 (ref:qwen2.5, ) and GPT-4o (ref:gpt4o, ) are employed as essential baselines to directly answer queries under the zero-shot setting. (2) Prompt Optimization Methods: RLPrompt (ref:rlprompt, ), BPO (ref:bpo, ), OPRO (ref:opro, ), and CFPO (ref:cfpo, ). For RLPrompt and BPO, we use Qwen2.5 as the base prompt generator, while for LLM-based methods such as OPRO and CFPO, we use GPT-4o as the prompt optimizer. (3) Fine-Tuning Methods: SFT, PPO (ref:ppo, ), DPO (ref:dpo, ), and KTO (ref:kto, ). For RL-based methods (PPO, DPO, and KTO), we first perform SFT as a warm-up stage and then construct alignment data following SPIN (ref:spin, ). (4) Hybrid Methods: BetterTogether (ref:better_together, ) is a hybrid method that iteratively optimizes the prompts and weights. However, it does not incorporate the parameter generation network and knowledge-sharing mechanism. As mentioned in the Section 3.2, we optimize our model with two strategies: In MetaTuner-I, we adopt a sequential optimization strategy, where the two terms in Equation 8 are optimized alternatively. In MetaTuner-J, we optimize Equation 8 in a unified manner.
Implementation details. We use Qwen2.5-7B and Qwen2.5-3B as the prompt generator $\mathcal{G}$ and the downstream actor model $\mathcal{M}$. Prior to joint training, we initiate an SFT stage to warm up $\mathcal{G}$ and $\mathcal{M}$ respectively. The training is carried out through Llama-Factory (ref:llamafactory, ), with the epoch set to $1$, batch size set to $64$, and the learning rate tuned in {1ee-3, 5ee-4, 1ee-4, 5ee-5}. Details of the stage can be found in the Appendix D.2. We then search the depth $k$ of $\phi_{s}$ in $\{K/4,K/2,3K/4,K\}$ to determine the model structure, where $K$ denotes the number of decoder layers in $\mathcal{G}$. In the training process, we search $r_{\mathcal{M}}$ in {4, 8, 16, 32}, $\alpha$ in {0.1, 0.5, 0.9}, the temperature $t$ for online rollouts in {0, 0.3, 0.5, 0.7, 0.9}, and the learning rate in {1ee-5, 5ee-6, 1ee-6, 5ee-7}. Additionally, we introduce a hyper-parameter $\lambda$ to scale the generated LoRA before adding it to the original model weights, and search $\lambda$ over {0.01, 0.05, 0.1, 0.5}. During the inference stage, we employ greedy decoding to ensure reproducibility. All the experiments are conducted on eight NVIDIA A100-80G GPUs.
Table 1: Comparing MetaTuner with baselines. We highlight the best and second-best results.
Qwen2.5-7B Qwen2.5-3B
Methods MATH GSM8K HotpotQA CosmosQA MATH GSM8K HotpotQA CosmosQA
Vanilla Methods
Qwen2.5 18.44 51.63 19.85 36.80 17.11 49.36 3.95 25.00
GPT-4o 18.89 68.46 54.15 76.50 18.89 68.46 54.15 76.50
Prompt Optimization Methods
RLPrompt 31.33 53.15 43.00 81.20 22.67 45.94 33.15 67.15
BPO 32.67 58.00 43.90 82.05 26.44 53.90 40.45 72.90
OPRO 22.00 75.06 25.55 69.10 21.00 62.62 10.55 69.35
CFPO 27.11 55.64 43.45 84.80 17.78 55.27 20.05 64.20
Fine-Tuning Methods
SFT 41.33 61.41 43.20 82.65 29.33 56.18 49.00 52.05
PPO 41.78 62.02 45.85 84.10 34.22 55.95 51.30 52.25
DPO 43.78 63.68 44.70 87.90 36.59 60.42 41.50 61.20
KTO 41.78 62.47 45.05 85.00 33.78 57.70 46.90 58.70
Hybrid Methods
BetterTogether 41.56 67.93 52.30 89.80 35.78 66.80 41.10 86.60
MetaTuner-I 48.22 78.54 55.75 92.15 40.89 73.46 59.05 86.55
MetaTuner-J 48.67 78.92 54.56 92.25 41.33 73.08 58.85 87.15
4.2 Overall Performance
The overall comparison is reported in Table 1, where several observations can be drawn: (1) The hybrid optimization methods consistently outperform standalone prompt optimization and fine-tuning methods, highlighting the complementary strengths of prompt and parameter optimization. (2) Compared to BetterTogether, two variants of MetaTuner achieve an average relative improvement of 10.15% and 17.08% under the 7B and 3B backbone settings, respectively. This demonstrates that our design, featuring a shared meta encoder and a supervised regularization loss, captures the deeper synergies between prompts and parameters. (3) On most tasks, the MetaTuner-J variant yields slightly stronger results than the MetaTuner-I, suggesting that simultaneous optimization encourages consistent alignment between prompts and parameters. While on HotpotQA, the MetaTuner-I variant achieves better performance, with an average absolute improvement of 0.70% across the two backbone settings. This indicates that alternating updates may mitigate gradient interference between the two branches, leading to more stable training, which is particularly beneficial for challenging optimization scenarios. For subsequent analyses, we primarily focus on the MetaTuner-J variant.
4.3 Ablation Studies
In the above section, we evaluate our model as a whole. In this section, we would like to see how different components in our model contribute to the final performance. In specific, we compare our final model with its three variants: In MetaTuner (w/o F), we remove the fine-tuning component, that is, omitting the first term in Equation 8. In MetaTuner (w/o P), we remove the prompt optimization component, that is, omitting the second term in Equation 8. In MetaTuner (w/o S), we do not share the parameters between the prompt and parameter decoder, that is, we set $k=0$ and equip $\mathcal{F}$ with an independent $\phi_{s}^{\prime}$ with the same structure as $\mathcal{G}$. From the results in Table 2, we can see: Both terms in our supervised regularization loss play essential roles, since removing either of them leads to an average absolute accuracy drop of approximately 0.99% and 1.12%, respectively. These findings indicate that, rather than relying on the training of a single branch to search for the optimal prompt–parameter combination, jointly training both branches yields superior performance. Besides, instead of a simple combination of prompt optimization and fine-tuning, our framework enables two components to share common parameters in $\phi_{s}$ for mutual enhancement. The inferior performance of MetaTuner (w/o S) compared to the full model demonstrates the effectiveness of this design.
Table 2: Ablation Studies. We highlight the most and the second-most effective modules.
Qwen2.5-7B Qwen2.5-3B
Methods MATH GSM8K HotpotQA CosmosQA MATH GSM8K HotpotQA CosmosQA
MetaTuner (w/o F) 48.00 77.79 54.05 91.10 40.44 72.18 56.95 86.40
MetaTuner (w/o P) 46.22 78.54 53.90 91.00 39.78 72.25 57.65 86.55
MetaTuner (w/o S) 46.67 77.86 53.65 91.50 40.22 72.10 57.55 86.65
MetaTuner 48.67 78.92 54.56 92.25 41.33 73.08 58.85 87.15
4.4 Effectiveness of the Shared-Private Co-Optimization Structure
Figure 2: Performance with different proportions of shared decoder layers.
As demonstrated in the above experiments, parameter sharing between the fine-tuning and prompt optimization components is important for performance improvement. To further study the relation between the amount of shared parameters and the model performance, in this section, we examine how performance varies with different values of the sharing depth $k$ (introduced in Section 3.4), where $k$ is set to $\{K/4,K/2,3K/4,K\}$ and $K$ denotes the maximum number of decoder layers in $\mathcal{G}$. From the results in Figure 2, we can observe that for the 7B model, a smaller sharing ratio (K/4) yields the best results, whereas for the 3B model, a larger sharing ratio (3K/4) performs better. This highlights the trade-off between the shared and private parameters: for the larger 7B model, its representation capacity is sufficiently strong, and excessive sharing may restrict the specialization of the prompt and parameter branches. Preserving more private layers therefore allows the model to better exploit its potential. In contrast, for the smaller 3B model, a higher sharing ratio strengthens representational consistency between the two branches, thereby improving training stability and downstream performance.
4.5 Effectiveness of the Supervised Regularization Loss
In our framework, a key design for addressing the mixed discrete–continuous optimization problem is the incorporation of the supervised regularizer in the objective (8). To evaluate its effectiveness, we conduct comprehensive experiments based on four research questions: (1) Is our method superior to other well-known differentiable strategies such as Gumbel-Softmax? (2) In selecting the dataset $D_{2}$, should it be derived from the optimized model itself or from a stronger model such as GPT-4o? (3) How does the number of rollout samples affect performance? (4) How does the update frequency of $\phi_{p}$ influence the final performance? In the following, we report and analyze the experiment results.
Figure 3: Comparison results between our method with Gumbel-Softmax.
Comparing with other differentiable strategies. In this section, we compare our method with Gumbel-Softmax, where the categorical prompt token sampling is replaced with a Gumbel-Softmax distribution to enable differentiable sampling. As shown in Figure 3, our method significantly outperforms Gumbel-Softmax. The main reason is that Gumbel-Softmax relies on continuous relaxation to approximate the discrete sampling. Although this ensures differentiability in theory, such softening operations introduce gradient bias, so that the generated prompts may deviate semantically from the truly optimal discrete solutions. In contrast, our methods directly optimize over the discrete action space using task rewards, thereby avoiding the potential approximation errors.
Figure 4: Comparison results between different methods for generating $D_{2}$.
Comparison between different generation methods of $D_{2}$. We compare the performances by generating $D_{2}$ using GPT-4o or the optimized model itself. As shown in Figure 4, while the GPT-4o method improves training efficiency, our model achieves better performance. The key reasons are threefold: first, our model ensures distribution alignment by generating data directly from the current policy, thereby reducing distribution shift between training and inference; second, it provides more targeted feedback, as rewards are based on the model’s own outputs and directly address its weaknesses, unlike GPT-4o samples that may overlook error patterns; third, our model enables stronger exploration, allowing the policy to continuously discover new prompts and iteratively refine itself through task feedback, whereas the GPT-4o data remain static and limit long-term improvement.
Influence of the number of rollout samples. When building $D_{2}$, an important factor is how many rollout samples should be generated for each query. In this section, we study the relation between this factor and the model performance. The results in Figure 6 show that as the number of rollout samples increases, the performance of both model sizes declines. On one hand, over-exploration may cause the model to frequently adjust its strategy, and these adjustments may not lead to short-term performance improvements, potentially interfering with valuable information the model has already learned. On the other hand, excessive exploration can cause overfitting, making the model overly sensitive to noise or irrelevant features in the training data, which ultimately degrades the downstream performance. Therefore, the number of rollout samples should be carefully adjusted according to the specific task and model characteristics to avoid the negative impact of excessive exploration.
Figure 5: Influence of the rollout number.
Figure 6: Influence of the update frequency.
Influence of the update frequency of $\phi_{p}$. In the first term of our objective (8), the private parameters $\phi_{p}$ for prompt generation is updated in an asynchronous manner to ensure that the objective is fully differentiable. Here, we investigate the influence of different update frequencies and report the results in Figure 6. On the horizontal axis, the values indicate how many training steps elapse before $\phi_{p}$ used for rollout sampling is replaced with the most recently trained one. The results on both model sizes consistently show that sparser updates lead to poorer performance, since frequent updates allow the sampled prompts to be closely aligned with the current actor model, thereby providing more effective optimization signals. Moreover, we observe that the 3B model exhibits a greater performance drop, which suggests that smaller models are more sensitive to the timeliness of generated prompts, whereas the 7B model, with its stronger representational capacity, is more resilient to such lag effects.
4.6 Generalization Analysis
Figure 7: Generalization experiment.
In the above sections, we evaluate our model by training and testing on the same dataset. To further study the generalization capability of our method, we conduct out-of-distribution experiments by training the model on MATH, HotpotQA, and CosmosQA, and evaluate it on GSM8K. For comparison, we select DPO, BPO, and BetterTogether as the baselines. From the results in Figure 7, we can see: MetaTuner outperforms all other methods, confirming the advantage of joint optimization of prompts and parameters in enhancing the generalization ability. Compared to BetterTogether, MetaTuner achieves better synergistic optimization, thereby adapting more effectively to complex and unseen tasks. DPO, as a fine-tuning method, although effective in parameter adjustment, lacks support in prompt selection, while BPO, which relies solely on prompt optimization and fails to fully leverage the model’s capacity, both lead to poorer generalization performance.
5 Related Work
Fine-tuning. Fine-tuning is a standard approach for adapting pre-trained language models to specific tasks. Recent advances in parameter-efficient fine-tuning (PEFT) scale this to LLMs by training small plug-in modules under a frozen backbone. Representative methods include Adapter Tuning (ref:adapter_tuning, ), which inserts lightweight adapter layers, and LoRA (ref:lora, ), which applies low-rank updates to weight matrices. Building on these, recent works propose various post-training strategies (ref:kto, ; ref:online_dpo, ; ref:bco, ; ref:cpo, ; ref:opro, ). For instance, RLHF (ref:rlhf, ) incorporates human feedback to improve alignment, while DPO (ref:dpo, ) streamlines preference modeling through contrastive learning.
Prompt optimization. Prompt optimization has become a research hotspot with the rise of in-context learning. Early works focus on continuous prompt tuning (ref:prefix_tuning, ; ref:google_soft_prompt, ; ref:p_tuning, ; ref:p_tuning_v2, ; ref:prompt_tuning, ), where learnable vectors are attached to the input. To mitigate the interpretability challenge (ref:hard_better_1, ), recent studies turn to discrete prompt optimization, which generates prompts in natural language. Existing approaches fall into two categories. The first employs LLMs as prompt optimizers (ref:hmaw, ), typically in a feedback-driven manner, via strategies such as reflection-based refinement (ref:opro, ; ref:llm_dual_phase_po, ; ref:grad_sum, ) or Monte Carlo Tree Search (MCTS) (ref:prompt_agent, ; ref:protegi, ; ref:cfpo, ). The second (ref:qpo, ; ref:uniprompt, ; ref:rlprompt_improve, ; ref:prompt_agent, ; ref:mapo, ) trains an auxiliary model to generate better prompts, most often with reinforcement learning, since prompts lack explicit ground truth and can be naturally guided by the task rewards.
Joint optimization. Recently, BetterTogether (ref:better_together, ) proposes to update prompts and model weights together to improve the task performance. In BetterTogether, prompt optimization and fine-tuning correspond to discrete and continuous optimization problems, respectively. During the prompt optimization phase, it samples candidate demonstration sets from the training data, evaluates them on a validation set to select the best few-shot prompt. In contrast, the fine-tuning phase is formulated as a continuous optimization problem, where supervised fine-tuning is applied to update the model parameters for better task adaptation. Different from BetterTogether, we first transform the discrete–continuous optimization into a fully continuous problem, then propose a knowledge sharing mechanism and a supervised regularization loss to enable effective end-to-end co-optimization.
6 Conclusion
In this paper, we propose MetaTuner, a novel framework that effectively combines prompt optimization and fine-tuning to enhance the performance of LLMs. We first introduce a prompt generator to transform the discrete–continuous optimization problem into a fully continuous one, and then design a dual-branch framework with a shared meta encoder to generate prompts and parameters simultaneously. We further propose a supervised regularization loss to address the non-differentiability challenge, thereby enabling stable end-to-end training of the framework. Extensive experiments on four benchmarks demonstrate the effectiveness and robustness of MetaTuner for joint optimization.
Reproducibility Statement
We take several steps to ensure the reproducibility of our work. Firstly, we detail the joint optimization objective and supervised regularization loss in Sections 3.2 and 3.3, and present the architectural details in Section 3.4. These descriptions are intended to make both the methodological design and implementation choices transparent. Secondly, we document the experimental setup in Section 4.1, including datasets, evaluation metrics, and training configurations, to enable the replication of the training and evaluation procedures. We further elaborate on data splits, the warm-up stage, and the meta-prompt design in the Appendix D.2 to facilitate replication. We believe these efforts provide the community with sufficient resources to reproduce our work and extend it in future research.
Ethics Statement
The development and deployment of frameworks like MetaTuner, which aim to enhance the performance of large language models through joint prompt-parameter optimization, raise several ethical considerations. First, as with any system based on large-scale language models, there is a risk of inheriting or amplifying biases present in the training data. If not carefully addressed, the improved adaptability of MetaTuner could inadvertently reinforce harmful stereotypes or generate misleading information. Second, the computational resources required to train and fine-tune large language models contribute to environmental concerns related to energy consumption and carbon footprint. Future work should explore new ways to further reduce the resource demands.
Appendix A LLM Usage
The usage of LLMs in this work includes two aspects: (1) Writing assistance. We use LLMs to improve the fluency and clarity of certain parts of the manuscript. All generated text is carefully reviewed, revised, and validated by the authors to ensure accuracy. (2) Dataset construction. We employ LLMs as auxiliary tools to generate candidate prompts as described in the Appendix D.2. After task-reward filtering, the collected data are used to improve the open-source smaller models’ ability to follow prompt-rewriting instructions. All scientific claims, methodological designs, implementations, analyses, and conclusions are original contributions of the authors. LLMs are used solely as supporting tools, and the authors take full responsibility for the integrity and accuracy of the work.
Appendix B Limitations
While MetaTuner presents a promising framework for jointly optimizing prompts and parameters, there are several limitations that point to potential directions for future research. First, the RL-based approach used to address the non-differentiability issue in prompt optimization relies heavily on task reward signals (e.g., 0/1 accuracy). Such signals can be overly sparse in certain tasks, which may hinder optimization efficiency. Future work could explore more fine-grained reward designs, such as model uncertainty, to provide richer and more informative optimization signals. Second, MetaTuner generates a dedicated set of prompt and parameters for each query, which may face computational challenges in resource-constrained environments. A promising direction is to integrate the Mixture of Experts (MoE) paradigm: generate and maintain a fixed pool of expert LoRA modules, and dynamically route each query to the most suitable expert(s). This design could avoid generating a new LoRA for every input, thereby improving the efficiency and scalability.
Appendix C Additional Experiments
C.1 Case Study
We present a case from the MATH dataset in Figure 8 to illustrate the advantages of MetaTuner over existing methods. Compared with pure prompt optimization approaches such as BPO, which can produce step-by-step reasoning but often generate overly lengthy intermediate steps and occasionally commit calculation errors that lead to incorrect answers, MetaTuner produces reasoning traces that are both accurate and concise while strictly adhering to the required output format. This improvement stems from its joint optimization of prompts and parameters, which allows the model to better adapt to downstream tasks. In contrast, when compared with traditional fine-tuning approaches such as DPO, MetaTuner also demonstrates clear advantages. Although DPO can arrive at the correct numerical solution, it fails to comply with formatting requirements and tends to generate relatively rigid and less context-aware reasoning. MetaTuner, by dynamically generating high-quality prompts that activate relevant internal knowledge while simultaneously refining model parameters, effectively combines the strengths of both prompt optimization and fine-tuning. As a result, MetaTuner achieves more stable, flexible, and diverse reasoning patterns, demonstrating both correctness and expressiveness.
Figure 8: Case study from the MATH dataset.
C.2 Efficiency Analysis
During the inference stage, for a given query $x$, the MetaTuner framework needs to generate a set of prompts and parameters to complete the downstream task. Unlike the standard text generation task, there are additional costs associated with the LoRA parameter generation. In practice, we generate parameters for the o_proj part of the decoder layers in the downstream actor model $\mathcal{M}$. As mentioned in Section 3.4, each LoRA parameter is generated by a hyper-network $\phi_{q}=\{W_{d}^{b}\in\mathbb{R}^{d_{\mathcal{M}}\times l},W_{u}^{b}\in\mathbb{R}^{h_{\mathcal{G}}\times r_{\mathcal{M}}},W_{d}^{a}\in\mathbb{R}^{r_{\mathcal{M}}\times l},W_{u}^{a}\in\mathbb{R}^{h_{\mathcal{G}}\times k_{\mathcal{M}}}\}$. Therefore, depending on the task scenario and model choice, the additional cost varies with sequence length $l$, the hidden size $h_{\mathcal{G}}$ of the prompt generator, the LoRA rank $r_{\mathcal{M}}$ of the generated parameters, and the number of decoder layers $K$ in the downstream actor model. However, compared to the parameter quantity of the LLM itself, the hyper-network is relatively small, so it does not incur significant computational overhead. In Table 3, we present the inference costs on a single NVIDIA A100 80GB GPU, measuring the average time cost per query for parameter generation across different datasets and base models.
Table 3: Average inference cost of parameter generation per query (second).
Base Model MATH GSM8K HotpotQA CosmosQA Avg
Qwen2.5-7B 0.13 0.11 0.02 0.30 0.14
Qwen2.5-3B 0.49 0.53 0.03 0.12 0.29
Appendix D Implementation Details
D.1 Preliminary Experiment
We present the prompts used in the right subfigure of the preliminary experiment below. For constructing the training data, the prompt $p$ is concatenated in front of the query $x$ to form the model input, while the ground truth $y$ serves as the expected output. The supervised fine-tuning process is conducted using Llama-Factory (ref:llamafactory, ), with the epoch set to 1, the batch size set to 64, and the learning rate tuned over the values {5ee-4, 1ee-4}.
Prompts for Supervised Fine-Tuning on HotpotQA. ## Prompt_0 From the given text, define or describe a particular term, concept, or idea mentioned. Your response should be based solely on how the term is presented in the passage. ## Prompt_1 Based on the passage, identify the character or person associated with the provided description. Provide the correct name or identifier, as supported by the text. ## Prompt_2 Use the passage to infer the implied meaning or consequence of a particular event, decision, or fact. Provide the most logical answer based on the information available. ## Prompt_3 Using the passage provided, identify the location or place associated with the subject matter. Make sure the answer is clearly supported by information found in the text. ## Prompt_4 Read the provided text and identify specific facts related to a particular subject. Answer the question using only information found in the passage. ## Prompt_5 Compare the details given in the passage and answer the question by identifying similarities or differences between two subjects or pieces of information. ## Prompt_6 From the given text, determine the timeline of events or specific dates related to a topic. Answer the question with precise chronological details from the passage. ## Prompt_7 Retrieve a specific detail or piece of information from the passage to answer a question. Make sure the answer directly aligns with what is stated in the text. ## Prompt_8 Given the passage, answer a comparative question by identifying similarities or differences between two concepts, characters, or events described in the text. ## Prompt_9 Identify the relationship between two subjects or concepts described in the passage. Answer the question by focusing on how these two entities are connected or related. ## Prompt_10 From the passage, answer the question by explaining the situation or problem presented in the text. Provide a clear response that addresses the core issue or situation described. ## Prompt_11 Read the passage and summarize the key points that answer the question. Focus on providing a concise, clear response that captures the main elements of the text. ## Prompt_12 Given a passage, answer the question by making an inference based on the information provided. Your answer should logically follow from the details in the text, even if it’s not explicitly stated. ## Prompt_13 From the passage, answer the question by providing a summary of the main events or points described. Your answer should focus on the key takeaways and provide a concise overview of what is happening. ## Prompt_14 Based on the passage, explain a process or sequence of events described in the text. Provide the steps or order of occurrences in your response. ## Prompt_15 From the passage, answer the question by considering the broader context in which the information is provided. Your answer should reflect an understanding of the context and its relevance to the question. ## Prompt_16 Based on the passage, answer the question by extracting the most important details that would influence a decision or action. Your answer should be based on the most relevant information provided in the text. ## Prompt_17 From the provided passage, answer the question by identifying the key detail or piece of information that directly addresses the question. ## Prompt_18 Given a passage, extract the relevant information that answers a specific question. Ensure that your response is directly supported by details in the passage. ## Prompt_19 Based on the passage, answer the question by identifying the character or event described. Provide the answer that best fits the situation outlined in the text.
D.2 Implementation of MetaTuner
Data splition. We split each dataset into training, validation, and test sets for model training and evaluation. For MATH, we use the 12K instances in the training set as specified in (ref:math_data_split, ), and randomly divide the remaining 500 instances into 50 for validation and 450 for testing. For GSM8K, the official test set contains 1,319 instances, so we allocate 473 instances for validation and the remaining 7K instances for training. For HotpotQA, we sample 20,000 instances from the official training set and randomly select 500 for validation and 2,000 for testing from the official dev-set-distractor. For CosmosQA, we use the entire 25,262 instances in the training set and randomly sample 500 and 2,000 instances from the test set for validation and testing, respectively.
Details of the warm-up stage. For the warm-up of $\mathcal{M}$, we use the queries $x$ from the original dataset as model inputs and the corresponding ground truth $y$ as model outputs for supervised fine-tuning. For the warm-up of $\mathcal{G}$, since there is no ground truth prompt, we construct the supervised training data with rejection sampling. Specifically, we first utilize GPT-4o to generate expert prompts $\mathcal{D}_{\text{expert}}=\{\hat{p}_{i}\}_{i=1}^{N}$, and then use the task reward $\mathcal{R}$ to filter out positive examples. The positive samples are then collected as $\mathcal{D}_{\text{PO}}=\{(x_{i},\hat{p}_{i})|\mathcal{R}(\mathcal{M}^{*}(\hat{p}_{i},x_{i}))=1\}_{i=1}^{N}$ for the supervised fine-tuning of prompt generation. Here, $\mathcal{M}^{*}$ and $\mathcal{G}^{*}$ represent the model after warm-up. Naturally, the pre-warmed model can form a pipeline $\hat{y_{i}}=\mathcal{M}^{*}(\mathcal{G}^{*}(\tilde{p},x_{i}),x_{i})$ to complete downstream tasks. We report the results in Table 4.
Table 4: Performance after the warm-up stage.
Base Model MATH GSM8K HotpotQA CosmosQA
Qwen2.5-7B 45.33 77.48 52.20 90.05
Qwen2.5-3B 35.33 71.27 53.95 84.00
Construction of the meta prompts. In this paper, since we adopt a prompt-rewrite paradigm for prompt generation, each meta prompt contains a query and an initial prompt, as well as an instruction to describe the prompt rewrite task. We follow the guidance of (ref:llm_dual_phase_po, ) to construct a base version of the prompt for each task as the initial prompt, which only includes a description of the task and the required answer format to facilitate answer parsing. For GSM8K, since the base model already produces highly consistent answer formats, we omit the answer format requirement in the initial prompt. The meta prompts on four datasets are presented below.
[…]
Meta prompt for GSM8K dataset. Given a query and an initial prompt, please add more detailed guidance to the initial prompt to help the model better respond to the question. The generated prompt will be concatenated in front of the query to assist the model in answering questions. During the rewriting process, ensure that the new prompt includes the information from the original prompt, such as the task description and the answer format requirements. The generated new prompt should not exceed 100 words. Query: {query} Initial Prompt: Given a complex mathematical problem, solve it step by step using a chain-of-thought approach. New Prompt:
Meta prompt for HotpotQA dataset. Given a query and an initial prompt, please add more detailed guidance to the initial prompt to help the model better respond to the question. The generated prompt will be concatenated in front of the query to assist the model in answering questions. During the rewriting process, ensure that the new prompt includes the information from the original prompt, such as the task description and the answer format requirements. The generated new prompt should not exceed 100 words. Query: {query} Initial Prompt: Use the relevant information provided to answer the multi-hop reasoning question. Please provide the most accurate and concise answer to the question based on the information provided. The answer should be directly from the dataset, with no extra details or explanations. For yes-no questions, respond with either ’yes’ or ’no.’ Avoid long or complex sentences. New Prompt:
Meta prompt for CosmosQA dataset. Given a query and an initial prompt, please add more detailed guidance to the initial prompt to help the model better respond to the question. The generated prompt will be concatenated in front of the query to assist the model in answering questions. During the rewriting process, ensure that the new prompt includes the information from the original prompt, such as the task description and the answer format requirements. The generated new prompt should not exceed 100 words. Query: {query} Initial Prompt: Based on the provided context, select the most relevant and accurate option as the correct answer to the question. The answer should be directly related to the context, and must focus on the key details or events described. New Prompt: